# The Hidden Power of Normalization: Exponential Capacity Control in Deep Neural Networks

### v1

@@ A.2 Properties of Layer norm
$\leq$ $\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}\left\|({\bm{I}}-\frac{{\bm{y}}^{T}{\bm{y}}}{\|{\bm{y}}\|_{2}^{2}+n\epsilon})\right\|\left(1-\frac{1}{n}\right)$ (27)
Consider matrix ${\bm{A}}={\bm{I}}-\frac{{\bm{y}}^{T}{\bm{y}}}{\|{\bm{y}}\|_{2}^{2}+n\epsilon}$. Its element at row $i$ and column $j$ is $A_{ij}=\delta_{ij}-\frac{y_{i}y_{j}}{\|{\bm{y}}\|_{2}^{2}+n\epsilon}$, where $\delta_{ij}=\delta({i=j})$ denotes the indicator function. It is easy to show that $\|{\bm{y}}\|_{2}^{2}+n\epsilon-y_{i}y_{j}\geq 0$ for any $(i,j)$, and hence $|{y_{i}y_{j}}|\leq{\|{\bm{y}}\|_{2}^{2}+n\epsilon}$. This suggest $|A_{ij}|\leq 1$ and $\|{\bm{A}}\|\leq 1$. Combining this with (27) shows $\left\|\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|\leq\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}\left(1-\frac{1}{n}\right)$. Therefore,
@@
$\leq$ $\sup_{{\bm{x}}\in\mathbb{R}^{n}}\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}\left(1-\frac{1}{n}\right)$ (29)
$\leq$ $\sup_{\sigma_{n}}\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}\left(1-\frac{1}{n}\right)$ (30)
$\leq$ $\left(1-\frac{1}{n}\right)\epsilon^{-0.5}$ (31)
which completes the first statement.
For the second statement, the assumption about existence of $\sigma$ implies that $\left\|\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|\leq\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}\left(1-\frac{1}{n}\right)\leq\frac{1}{\sigma}\left(1-\frac{1}{n}\right)$ for any $\bm{x}$}. As a consequence, $\|{\mathsf{LN}}\|_{Lip}=\sup_{{\bm{x}}}\left\|\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|\leq\frac{1}{\sigma}\left(1-\frac{1}{n}\right)$, completing the proof.
@@ A.3 Properties of Group norm
Since the groups are disjoint and $\|\cdot\|$ is the max norm, we have
$\left\|\frac{\partial{\mathsf{GN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|$ $\leq$ $\max_{1\leq k\leq C}\left\|\frac{\partial{\mathsf{GN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}_{k}}\right\|$ (33)
Remember that for each group ${\bm{S}}_{k}$, the $\mathsf{GN}$} operator is in fact $\mathsf{LN}$}. Hence, $\frac{\partial{\mathsf{GN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}_{k}}=\frac{\partial{\mathsf{LN}}({\bm{x}}_{k},\epsilon)}{\partial{\bm{x}}_{k}}$. The proof of Lemma 2 implies $\left\|\frac{\partial{\mathsf{LN}}({\bm{x}}_{k},\epsilon)}{\partial{\bm{x}}_{k}}\right\|\leq\frac{1}{\sqrt{\sigma_{k}^{2}({\bm{x}})+\epsilon}}\left(1-\frac{1}{|{\bm{S}}_{k}|}\right)$. Combining this with (33) will arrive at
$\left\|\frac{\partial{\mathsf{GN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|\leq\max_{1\leq k\leq C}\frac{1}{\sqrt{\sigma_{k}^{2}({\bm{x}})+\epsilon}}\left(1-\frac{1}{|{\bm{S}}_{k}|}\right)$
@@
$\sup_{{\bm{x}}\in\mathbb{R}^{n}}\left\|\frac{\partial{\mathsf{GN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|$ $\leq$ $\sup_{{\bm{x}}\in\mathbb{R}^{n}}\max_{1\leq k\leq C}\frac{1}{\sqrt{\sigma_{k}^{2}({\bm{x}})+\epsilon}}\left(1-\frac{1}{|{\bm{S}}_{k}|}\right)$ (34)
$\leq$ $\max_{1\leq k\leq C}\frac{1}{\sqrt{\sigma_{k}^{2}+\epsilon}}\left(1-\frac{1}{|{\bm{S}}_{k}|}\right)$ (35)
$=$ $\|{\bm{v}}\|$ (36)
@@ Appendix B DNNs with normalizations
$\leq$ $\|g_{K}\|_{Lip}\|{\mathsf{NO}}_{K}\|_{Lip}\|{\bm{W}}_{K}\|\|{\bm{h}}_{K-1},{\bm{x}}\|_{Lip}$ (43)
$\leq$ $\|{\bm{W}}_{K}\|\|{\mathsf{NO}}_{K}\|_{Lip}\|{\bm{h}}_{K-1},{\bm{x}}\|_{Lip}$ (44)
$\cdots$ (45)
$\leq$ $\prod_{k=1}^{K}\|{\bm{W}}_{k}\|\|{\mathsf{NO}}_{k}\|_{Lip}$ (46)
@@
Consider a model ${\bm{h}}_{no}={\bm{h}}_{K}=g_{K}({\mathsf{NO}}_{K}({\bm{y}}_{K}))$. A basic property of Lipschitz continuity for composite functions suggests that $\left\|{\bm{h}}_{no},{\bm{y}}_{K}\right\|_{Lip}\leq\|g_{K}\|_{Lip}\|{\mathsf{NO}}_{K}\|_{Lip}\leq\|{\mathsf{NO}}_{K}\|_{Lip}$.
@@
$\left\|{\bm{h}}_{no},{\bm{y}}_{i}\right\|_{Lip}$ $=$ $\left\|{\bm{h}}_{K},{\bm{y}}_{i}\right\|_{Lip}$ (47)
$\leq$ $\|g_{K}\|_{Lip}\|{\mathsf{NO}}_{K}\|_{Lip}\|{\bm{W}}_{K}{\bm{h}}_{K-1},{\bm{y}}_{i}\|_{Lip}$ (48)
$\leq$ $\|g_{K}\|_{Lip}\|{\mathsf{NO}}_{K}\|_{Lip}\|{\bm{W}}_{K}\|\|{\bm{h}}_{K-1},{\bm{y}}_{i}\|_{Lip}$ (49)
$\leq$ $\|{\bm{W}}_{K}\|\|{\mathsf{NO}}_{K}\|_{Lip}\|{\bm{h}}_{K-1},{\bm{y}}_{i}\|_{Lip}$ (50)
$\cdots$ (51)
$\leq$ $(\prod_{k=i+2}^{K}\|{\bm{W}}_{k}\|\|{\mathsf{NO}}_{k}\|_{Lip})\|{\bm{h}}_{i+1},{\bm{y}}_{i}\|_{Lip}$ (52)
$\leq$ $(\prod_{k=i+2}^{K}\|{\bm{W}}_{k}\|\|{\mathsf{NO}}_{k}\|_{Lip})\|{\mathsf{NO}}_{i+1}\|_{Lip}\|{\bm{W}}_{i+1}\|\|{\bm{h}}_{i},{\bm{y}}_{i}\|_{Lip}$ (53)
$\leq$ $(\prod_{k=i+1}^{K}\|{\bm{W}}_{k}\|\|{\mathsf{NO}}_{k}\|_{Lip})\|{\mathsf{NO}}_{i}\|_{Lip}$ (54)
$\leq$ $(\prod_{k=i+1}^{K}\|{\bm{W}}_{k}\|)\prod_{k=i}^{K}\|{\mathsf{NO}}_{k}\|_{Lip}$ (55)
Similarly, we next consider the Lipschitz constant w.r.t. the weights. Observe that $\left\|{\bm{h}}_{no},{\bm{W}}_{i}\right\|_{Lip}\leq\left\|{\bm{h}}_{no},{\bm{y}}_{i}\right\|_{Lip}\|{\bm{y}}_{i},{\bm{W}}_{i}\|_{Lip}$. Since ${\bm{y}}_{i}={\bm{W}}_{i}{\bm{h}}_{i-1}$, remember that $\|{\bm{y}}_{i},{\bm{W}}_{i}\|_{Lip}=\sup_{{\bm{h}}_{i-1}}\|{\bm{h}}_{i-1}\|=A_{i-1}$. As a result, $\left\|{\bm{h}}_{no},{\bm{W}}_{i}\right\|_{Lip}\leq A_{i-1}\left\|{\bm{h}}_{no},{\bm{y}}_{i}\right\|_{Lip}$, completing the proof.
@@ Appendix C Unnormalized DNNs
$\|{\bm{h}}_{K},{\bm{y}}_{i}\|_{Lip}$ $=a_{K}a_{K-1}\cdots a_{i+1}\left\|ReLU\left({\bm{y}}_{i}^{(n_{i+1},...,n_{K})}\right),{\bm{y}}_{i}\right\|_{Lip}$ (74)
@@
$\|{\bm{h}}_{K},{\bm{W}}_{i}\|_{Lip}$ $=a_{K}a_{K-1}\cdots a_{i+1}\left\|ReLU\left({\bm{y}}_{i}^{(n_{i+1},...,n_{K})}\right),{\bm{W}}_{i}\right\|_{Lip}$ (77)
@@
Consider the member ${\bm{h}}^{*}({\bm{x}})=g({\bm{W}}{\bm{h}}_{K}({\bm{x}}))$ for ${\bm{W}}=c\mathbf{1}^{\top}$ and ${\bm{h}}_{K}$ whose the weight matrix at layer $k$ is ${\bm{W}}_{k}=a_{k}{\bm{I}}_{k}$ for all $k>i$, where ${\bm{I}}_{k}$ is defined in Equation (56). Following the same arguments as the proof of Theorem 13 before, we can obtain a similar result as Equation (73):
@@
$=\frac{1}{m}\sum_{{\bm{x}}\in{\bm{D}}}\frac{\partial(f\odot g)}{\partial u}\frac{\partial u}{\partial{\bm{h}}_{K}}\frac{\partial{\bm{h}}_{K}({\bm{x}})}{\partial{\bm{W}}_{it}}$ (84)
$=\frac{1}{m}\sum_{{\bm{x}}\in{\bm{D}}}\frac{\partial(f\odot g)}{\partial u}{\bm{W}}\frac{\partial{\bm{h}}_{K}({\bm{x}})}{\partial{\bm{W}}_{it}}$ (85)
@@
Consider vector ${\bm{v}}=ReLU\left(({\bm{W}}_{i}{\bm{h}}_{i-1})^{(n_{i+1},...,n_{K})}\right)$, which is the series of truncation/zero-padding to ${\bm{W}}_{i}{\bm{h}}_{i-1}$ and then followed by $ReLU$ operation. As a result the element at row $k$ is
$v_{k}=\begin{cases}{\bm{W}}_{ik}{\bm{h}}_{i-1}&\text{ if }k\leq\min\{n_{i+1},...,n_{K}\}\text{ and }{\bm{W}}_{ik}{\bm{h}}_{i-1}\geq 0\\ 0&\text{ otherwise}\end{cases}$ (86)
@@
$\frac{\partial L}{\partial{\bm{W}}_{it}}=\frac{ca_{K}\cdots a_{i+1}}{m}\sum\limits_{{\bm{x}}\in{\bm{D}}}\frac{\partial(f\odot g)}{\partial u}{\mathcal{I}}[{\bm{W}}_{it}{\bm{h}}_{i-1}({\bm{x}})\geq 0]({\bm{h}}_{i-1}({\bm{x}}))^{\top}$

### v2

@@ A.2 Properties of Layer norm
$\frac{\partial{\mathsf{LN}}}{\partial{\bm{y}}}=\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}\left({\bm{I}}-\frac{{\bm{y}}{\bm{y}}^{T}}{\|{\bm{y}}\|^{2}+n\epsilon}\right)$ (18)
Note that $\frac{\partial{\bm{y}}}{\partial{\bm{x}}}={\bm{I}}-\frac{1}{n}\bm{1}$. By the chain rule $\frac{\partial{\mathsf{LN}}}{\partial{\bm{x}}}=\frac{\partial{\mathsf{LN}}}{\partial{\bm{y}}}\frac{\partial{\bm{y}}}{\partial{\bm{x}}}$, we obtain:
$\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}=\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}\left({\bm{I}}-\frac{{\bm{y}}{\bm{y}}^{T}}{\|{\bm{y}}\|^{2}+n\epsilon}\right)\left({\bm{I}}-\frac{1}{n}\bm{1}\right)$
Proof [Proof of Lemma 2] From Lemma 23, the Jacobian of the Layer Normalization function is given by:
$\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}=\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}\left({\bm{I}}-\frac{{\bm{y}}{\bm{y}}^{T}}{\|{\bm{y}}\|^{2}+n\epsilon}\right)\left({\bm{I}}-\frac{1}{n}\bm{1}\right)$ (19)
where ${\bm{y}}=({\bm{I}}-\frac{1}{n}\bm{1}){\bm{x}}$ is the centered input vector. We seek to bound the spectral norm of this Jacobian.
Using the sub-multiplicative property of the spectral norm, we have:
$\left\|\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|$ $\leq$ $\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}\left\|{\bm{I}}-\frac{{\bm{y}}{\bm{y}}^{T}}{\|{\bm{y}}\|^{2}+n\epsilon}\right\|\left\|{\bm{I}}-\frac{1}{n}\bm{1}\right\|$
We analyze the two matrix terms separately:
  1. 1.
    Let ${\bm{P}}_{1}={\bm{I}}-\frac{1}{n}\bm{1}$. This is the centering matrix, which projects vectors onto the subspace orthogonal to the vector of all ones. As a projection matrix (and assuming $n>1$), its eigenvalues are either 0 or 1. Thus, its spectral norm is $\|{\bm{P}}_{1}\|=1$.
  2. 2.
    Let ${\bm{P}}_{2}={\bm{I}}-\frac{{\bm{y}}{\bm{y}}^{T}}{\|{\bm{y}}\|^{2}+n\epsilon}$.
    • If $\epsilon=0$, ${\bm{P}}_{2}$ is a standard projection matrix orthogonal to $\bm{y}$}, and $\|{\bm{P}}_{2}\|=1$.
    • If $\epsilon>0$, ${\bm{P}}_{2}$ scales the component of a vector parallel to $\bm{y}$} by a factor of $1-\frac{\|{\bm{y}}\|^{2}}{\|{\bm{y}}\|^{2}+n\epsilon}=\frac{n\epsilon}{\|{\bm{y}}\|^{2}+n\epsilon}\in(0,1)$, while leaving components orthogonal to $\bm{y}$} unchanged (eigenvalue 1). Thus, the maximum singular value is still 1, so $\|{\bm{P}}_{2}\|\leq 1$.
Substituting these bounds back into the inequality:
$\left\|\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|$ $\leq$ $\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}\cdot 1\cdot 1$ (20)
$=$ $\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}$ (21)
Finally, taking the supremum over all $\bm{x}$}:
@@
$\leq$ $\sup_{{\bm{x}}\in\mathbb{R}^{n}}\frac{1}{\sqrt{\sigma^{2}_{n}({\bm{x}})+\epsilon}}$ (23)
$=$ $\frac{1}{\sigma}$ (24)
which completes the proof.
@@ A.3 Properties of Group norm
$\left\|\frac{\partial{\mathsf{GN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|$ $\leq$ $\sum_{k=1}^{C}\left\|\frac{\partial{\mathsf{GN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}_{k}}\right\|$ (26)
Remember that for each group ${\bm{S}}_{k}$, the $\mathsf{GN}$} operator is in fact $\mathsf{LN}$}. Hence, $\frac{\partial{\mathsf{GN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}_{k}}=\frac{\partial{\mathsf{LN}}({\bm{x}}_{k},\epsilon)}{\partial{\bm{x}}_{k}}$. The proof of Lemma 2 implies $\left\|\frac{\partial{\mathsf{LN}}({\bm{x}}_{k},\epsilon)}{\partial{\bm{x}}_{k}}\right\|\leq\frac{1}{\sqrt{\sigma_{k}^{2}({\bm{x}})+\epsilon}}$. Combining this with (26) will arrive at
$\left\|\frac{\partial{\mathsf{GN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|\leq\sum_{k=1}^{C}\frac{1}{\sqrt{\sigma_{k}^{2}({\bm{x}})+\epsilon}}$
Therefore
$\sup_{{\bm{x}}\in\mathbb{R}^{n}}\left\|\frac{\partial{\mathsf{GN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|$ $\leq$ $\sup_{{\bm{x}}\in\mathbb{R}^{n}}\sum_{k=1}^{C}\frac{1}{\sqrt{\sigma_{k}^{2}({\bm{x}})+\epsilon}}$ (27)
$\leq$ $\sum_{k=1}^{C}\frac{1}{\sqrt{\sigma_{k}^{2}+\epsilon}}$ (28)
@@ Appendix B DNNs with normalizations
$\leq$ $\|g_{K}\|_{Lip}\|{\mathsf{NO}}_{K}\|_{Lip}\|{\bm{W}}_{K}\|\|{\bm{h}}_{K-1},{\bm{x}}\|_{Lip}$
$\leq$ $\|{\bm{W}}_{K}\|\|{\mathsf{NO}}_{K}\|_{Lip}\|{\bm{h}}_{K-1},{\bm{x}}\|_{Lip}$
$\cdots$
$\leq$ $\prod_{k=1}^{K}\|{\bm{W}}_{k}\|\|{\mathsf{NO}}_{k}\|_{Lip}$
@@
Consider a model ${\bm{h}}_{no}={\bm{h}}_{K}=g_{K}({\mathsf{NO}}_{K}({\bm{y}}_{K}))$. A basic property of Lipschitz continuity for composite functions suggests that
$\left\|{\bm{h}}_{no},{\bm{y}}_{K}\right\|_{Lip}\leq\|g_{K}\|_{Lip}\|{\mathsf{NO}}_{K}\|_{Lip}\leq\|{\mathsf{NO}}_{K}\|_{Lip}$.
@@
$\left\|{\bm{h}}_{no},{\bm{y}}_{i}\right\|_{Lip}=\left\|{\bm{h}}_{K},{\bm{y}}_{i}\right\|_{Lip}$
$\leq\|g_{K}\|_{Lip}\|{\mathsf{NO}}_{K}\|_{Lip}\|{\bm{W}}_{K}{\bm{h}}_{K-1},{\bm{y}}_{i}\|_{Lip}$
$\leq\|g_{K}\|_{Lip}\|{\mathsf{NO}}_{K}\|_{Lip}\|{\bm{W}}_{K}\|\|{\bm{h}}_{K-1},{\bm{y}}_{i}\|_{Lip}$
$\leq\|{\bm{W}}_{K}\|\|{\mathsf{NO}}_{K}\|_{Lip}\|{\bm{h}}_{K-1},{\bm{y}}_{i}\|_{Lip}$
$\cdots$
$\leq(\prod_{k=i+2}^{K}\|{\bm{W}}_{k}\|\|{\mathsf{NO}}_{k}\|_{Lip})\|{\bm{h}}_{i+1},{\bm{y}}_{i}\|_{Lip}$
$\leq(\prod_{k=i+2}^{K}\|{\bm{W}}_{k}\|\|{\mathsf{NO}}_{k}\|_{Lip})\|{\mathsf{NO}}_{i+1}\|_{Lip}\|{\bm{W}}_{i+1}\|\|{\bm{h}}_{i},{\bm{y}}_{i}\|_{Lip}$
$\leq(\prod_{k=i+1}^{K}\|{\bm{W}}_{k}\|\|{\mathsf{NO}}_{k}\|_{Lip})\|{\mathsf{NO}}_{i}\|_{Lip}$
$\leq(\prod_{k=i+1}^{K}\|{\bm{W}}_{k}\|)\prod_{k=i}^{K}\|{\mathsf{NO}}_{k}\|_{Lip}$
Similarly, we next consider the Lipschitz constant w.r.t. the weights. Note that $\left\|{\bm{h}}_{no},{\bm{W}}_{i}\right\|_{Lip}\leq\left\|{\bm{h}}_{no},{\bm{y}}_{i}\right\|_{Lip}\|{\bm{y}}_{i},{\bm{W}}_{i}\|_{Lip}$. Furthermore, $\|{\bm{y}}_{i},{\bm{W}}_{i}\|_{Lip}=\sup_{{\bm{h}}_{i-1}}\|{\bm{h}}_{i-1}\|=A_{i-1}$ due to ${\bm{y}}_{i}={\bm{W}}_{i}{\bm{h}}_{i-1}$. As a result, $\left\|{\bm{h}}_{no},{\bm{W}}_{i}\right\|_{Lip}\leq A_{i-1}\left\|{\bm{h}}_{no},{\bm{y}}_{i}\right\|_{Lip}$, completing the proof.
@@ Appendix C Unnormalized DNNs
$\|{\bm{h}}_{K},{\bm{y}}_{i}\|_{Lip}$
$=a_{K}a_{K-1}\cdots a_{i+1}\left\|ReLU\left({\bm{y}}_{i}^{(n_{i+1},...,n_{K})}\right),{\bm{y}}_{i}\right\|_{Lip}$ (53)
@@
$\|{\bm{h}}_{K},{\bm{W}}_{i}\|_{Lip}$
$=a_{K}a_{K-1}\cdots a_{i+1}\left\|ReLU\left({\bm{y}}_{i}^{(n_{i+1},...,n_{K})}\right),{\bm{W}}_{i}\right\|_{Lip}$ (56)
@@
Consider ${\bm{h}}^{*}({\bm{x}})=g({\bm{W}}{\bm{h}}_{K}({\bm{x}}))$ for ${\bm{W}}=c\mathbf{1}^{\top}$ and ${\bm{h}}_{K}$ whose the weight matrix at layer $k$ is ${\bm{W}}_{k}=a_{k}{\bm{I}}_{k}$ for all $k>i$, where ${\bm{I}}_{k}$ is defined in Equation (35). Following the same arguments as the proof of Theorem 13 before, we can obtain a similar result as Equation (52):
@@
$=\frac{1}{m}\sum_{{\bm{x}}\in{\bm{D}}}\frac{\partial(f\circ g)}{\partial u}\frac{\partial u}{\partial{\bm{h}}_{K}}\frac{\partial{\bm{h}}_{K}({\bm{x}})}{\partial{\bm{W}}_{it}}$ (63)
$=\frac{1}{m}\sum_{{\bm{x}}\in{\bm{D}}}\frac{\partial(f\circ g)}{\partial u}{\bm{W}}\frac{\partial{\bm{h}}_{K}({\bm{x}})}{\partial{\bm{W}}_{it}}$ (64)
@@
Consider vector ${\bm{v}}=ReLU\left(({\bm{W}}_{i}{\bm{h}}_{i-1})^{(n_{i+1},...,n_{K})}\right)$, which is the series of truncation/zero-padding to ${\bm{W}}_{i}{\bm{h}}_{i-1}$ and then followed by ReLUReLU operation. As a result the element at row $k$ is
$v_{k}=\begin{cases}{\bm{W}}_{ik}{\bm{h}}_{i-1}&\text{ if }k\leq\min\{n_{i+1},...,n_{K}\}\text{ and }{\bm{W}}_{ik}{\bm{h}}_{i-1}\geq 0\\ 0&\text{ otherwise}\end{cases}$
@@
$\frac{\partial L}{\partial{\bm{W}}_{it}}=\frac{ca_{K}\cdots a_{i+1}}{m}\sum\limits_{{\bm{x}}\in{\bm{D}}}\frac{\partial(f\circ g)}{\partial u}{\mathcal{I}}_{it}({\bm{h}}_{i-1}({\bm{x}}))^{\top}$